# SpikeSMOKE: Spiking Neural Networks for Monocular 3D Object Detection with Cross-Scale Gated Coding

### v2

@@ III-C Light-weight Residual
A novel light-weight residual block, as shown in Figure 3, is introduced to decrease model parameters and computational load, consequently accelerating its training speed, by integrating depth-wise separable convolutions and membrane shortcut method. Specifically, the depth-wise convolution is used to conduct a separate convolution operation for each input channel, which implies that redundant computations can be significantly reduced because it perform only one convolution per channel. Then, a point-wise convolution is applied to the output of depth-wise convolution. After that, another depth-wise convolution is applied. It should be noted that a membrane-based shortcut pathway is utilized to remain identity mapping.
[…]
For the SpikeSMOKE with light-weight residual block, the convolution kernel size $k$ is $3\times 3$, $C_{\text{in}}=2C_{\text{out}}$, $W_{\text{out}}=\frac{1}{2}W_{\text{in}}$, $H_{\text{out}}=\frac{1}{2}H_{\text{in}}$. Then, the ratio of the computational effort can be calculated by the Equation (8):
$\frac{Cin\cdot Wout\cdot Hout(3k^{2}+2Cin)}{k^{2}\cdot CinCout\cdot Wout\cdot Hout(1+4Cin)}=\frac{1}{2Cin}+\frac{1}{27}.$ (8)
@@
$\frac{k^{2}\cdot Cin+Cin\cdot Caut+k^{2}Cout}{k^{2}Cin\cdot Cout+k^{2}Cout\cdot Cout}=\frac{1}{2Cin}+\frac{1}{27}.$ (9)
Observably, compared to the pre-improvement phase, the proposed light-weight residual block diminishes computational and parametric quantities by approximately 27 times.
@@ IV-A Datasets
Where N denotes the set of recall levels that represent the exact intervals, this experiment takes $R_{11}=\{\frac{1}{10},\frac{2}{10},\frac{3}{10},\cdots,\\ 1\}$. $f_{inter}(n)$ denotes the precision obtained after interpolation calculation, instead of averaging the precision values, the maximum precision greater than or equal to the current recall is taken in calculating the precision. It is defined as the Equation (11):
@@
Methods
Parameters
Power
3D Object Detection Birds’ Eye View
@@ Methods
SMOKE-ANN*(0.5) 19.51 2.17E+11 29.16 25.22 21.47 32.52 29.59 25.86
SpikeSMOKE(0.5) 19.51 5.97E+10 20.87 19.72 16.89 27.64 22.47 21.65
SpikeSMOKE-CSGC(0.5) 19.56 6.04E+10 28.83 22.75 19.44 36.23 26.88 25.75
SMOKE-ANN(0.7)[34] 19.51 2.17E+11 14.76 12.85 11.5 19.99 15.61 15.28
SMOKE-ANN*(0.7) 19.51 2.17E+11 12.03 11.14 10.92 17.97 13.08 12.06
SpikeSMOKE(0.7) 19.51 5.97E+10 8.96 7.49 7.31 12.07 9.95 9.32
SpikeSMOKE-CSGC(0.7) 19.56 6.04E+10 11.78 10.69 10.48 15.67 13.68 11.83
(0.5/0.7) indicates that the metrics are evaluated by $AP|_{R_{11}}$ at the 0.5/0.7 IoU thresholds.
@@ IV-B Experiments Setup
We used random horizontal flip, random scale and shift as data augmentation aiming to increase the diversity of training data. In the network design, we set the number of groups for group normalization to 32 and for less than 32, we take 16. based on the analysis in literature [59], we set $\begin{bmatrix}\bar{h,}\bar{w,}\overline{l,}\end{bmatrix}^{\top}=[1.63,1.53,3.88]^{\top}and\begin{bmatrix}\mu_{z},\sigma_{z}\end{bmatrix}^{\top}=[28.01{\operatorname*{,16.32}}]^{\top}.$ The input image resolution size of the network is 1280x384, and after multiple downsampling, the size is reduced to 32 times of the original size. We trained 172 epoch using $4\times 4090$ GPUs, with a size of 4 per batch and a learning rate set to $1.25\times 10^{-4}$. The decay strategy for the learning rate was to reduce the learning rate to 10 times of the original at epoch 47 and 90. We used a threshold of 0.25 to filter the detection object.
@@
TABLE VI: Ablation experiments on the threshold of spike neurons.
@@ IV-C Main Results
We validate the performance on the validation and test sets of the KITTI dataset for 3D object detection and bird’s-eye view under three levels of Easy, Moderate and Hard. * denotes the experimental results obtained by us after configuring the environment and training based on the SMOKE. The 3d object detection of our proposed SpikeSMOKE with CSGC (SpikeSMOKE-CSGC) can reach 28.83/11.78 (Easy), 22.75/10.69 (Moderate), 19.44/10.48 (Hard), and for the BEV detection, they can reach 36.23/15.67 (Easy), 26.88/13.68 (Moderate), 25.75/11.83 (Hard), as shown in Table I, where the metrics is evaluated by $AP|_{R_{11}}$ at 0.5/0.7 IoU threshold. From the results of these experiments, it can be found that although our detection performance has a little gap compared with the SMOKE-ANN*, energy consumption has been significantly reduced. For example, we calculate the energy consumption of the hard category for the 0.7 IoU threshold and find that it can be reduced by 72.2%, while the detection performance is reduced by only 4%. The formula for calculating energy consumption is as follows:
$EC_{SNN}$ $=Synapsed_{activated}^{SNN}\times 0.9,$ (12)
$EC_{ANN}$ $=FLOPS^{ANN}\times 4.6,$
where 0.9 denotes the consumption per accumulator operation and 4.6 denotes the multiplier word operation.
@@
We validate the classification results of the proposed CSGC coding strategy on MS-ResNet18 on the classification dataset CIFAR-10/100, as shown in Table IV and Table V. We note that the accuracy rate of the MS-ResNet18 with CSGC coding on the CIFAR-10 is 1.06% higher than the direct coding method. On the CIFAR-100, the proposed method MS-ResNet18-SNN with CSGC coding can achieve 79.58% by 6 time steps, which is 3.17% higher than MS-ResNet18-SNN. Consequently, our proposed CSGC coding method is effective and generalized.
@@
Since the threshold of the spike neurons has a great impact on the detection performance of the model, we conduct some ablation experiments on 2D detection by different neuron thresholds in order to get the most suitable LIF neuron threshold. Based on the experiment results, we obtain that the best performance is obtained when the neuron threshold vth=0.75, which is used in this paper, as shown in Table VI.
@@
We provide visualization results of the monocular 3D object detection on the KITTI dataset as shown in Figure 4. Through these results, we can visually observe the model’s capability to accurately identify and locate 3D objects in complex road scenarios, thereby affirming the significant potential application of this technology in areas such as autonomous driving.
@@ V Conclusion
With the widespread use of 3D object detection in applications such as autonomous driving, the low energy consumption problem is getting more and more attention. As widely acknowledged, low-power consumption stands out as a key feature of brain-like SNNs, offering a potential new solution for energy-efficient 3D object detection. Based on this, we construct a novel SpikeSMOKE architecture for monocular 3D object detection. Since the discrete signaling characteristics of SNNs may result in information loss and restrict their capacity for feature representation, we propose a CSGC mechanism, inspired by the synaptic filtering process in biological neurons. Furthermore, we also propose a lightweight residual block to reduce the computational effort while maintaining the impulse computation paradigm. The experimental results on the KITTI dataset, show that SpikeSMOKE can achieve higher energy efficiency compared to SMOKE, e.g., 72.2% reduction in energy consumption on Hard category, while the detection performance drops by only 4%. Moreover, the experimental results also show that the CSGC-based SpikeSMOKE can achieve significant improvement over the baseline SpikeSMOKE. SpikeSMOKE-L (lightweight) can further reduce the amount of parameters by 3 times and computation by 10 times compared to SMOKE. In the CIFAR-10/100 classification task, the CSGC encoding strategy improves the correctness by 1.06% and 3.17%, respectively, validating its generality. Overall, the SpikeSMOKE architecture and CSGC mechanism can provide an efficient and feasible solution for low-power monocular 3D object detection.

### v3

@@ III-C Light-weight Residual
When SNN is applied to high-complexity tasks such as monocular 3D object detection, the residuals of standard convolution need to be repeatedly executed over multiple time steps, and the fully connected channel interaction can lead to a large amount of invalid computations, which goes against the original intention of energy efficiency of the SNN. Therefore, we introduce a novel light-weight residual block, as shown in Figure 3, is introduced to significantly reduce the number of model parameters and computational complexity by integrating depth-wise separable convolution technology and a membrane-based shortcut method, thereby enhancing the efficiency of model training. In terms of implementation details, the lightweight residual block first employs depth-wise convolution to perform independent convolution operations on each channel of the input data. Compared to traditional convolution, depth-wise convolution has the advantage of operating on each channel individually, thus avoiding redundant computations across channels and significantly improving computational efficiency. After completing the depth-wise convolution, we then apply point-wise convolution , with the purpose of integrating information from various channels and enhancing the expressive power of features. Then, we apply depth-wise convolution again to further extract and strengthen features. During this process, to maintain the flow of information and prevent the vanishing gradient problem, we introduce a shortcut path based on the membrane method
For regular membrane-based residual block, the computational volume $\alpha_{1}$ and number of parameters $\beta_{1}$ can be calculated by expression (6) and the expression (7) respectively:
$\alpha_{1}=k\cdot k\cdot C_{in}\cdot C_{out}\cdot W_{out}\cdot H_{out},$ (6)
$\beta_{1}=k\cdot k\cdot C_{in}\cdot C_{out}.$ (7)
When the above light-weight residual block is introduced, the computational volume $\alpha_{2}$ and parametric quantities $\beta_{2}$ can be calculated by the expression (8) and the expression (9) respectively:
$\alpha_{2}=k\cdot k\cdot C_{in}\cdot W_{out}\cdot H_{out}+C_{in}\cdot C_{out}\cdot W_{out}\cdot H_{out},$ (8)
$\beta_{2}=k\cdot k\cdot C_{in}+C_{in}\cdot C_{out}.$ (9)
For the SpikeSMOKE with light-weight residual block, the convolution kernel size $k$ is $3\times 3$, $C_{\text{out}}=2C_{\text{in}}$, $W_{\text{out}}=\frac{1}{2}W_{\text{in}}$, $H_{\text{out}}=\frac{1}{2}H_{\text{in}}$. Then, the ratio of the computational effort can be calculated by the Equation (10):
$\frac{C_{in}\cdot W_{out}\cdot H_{out}(3k^{2}+2C_{in})}{3k^{2}\cdot C_{in}\cdot C_{out}\cdot W_{out}\cdot H_{out}}=\frac{1}{2C_{in}}+\frac{1}{27}.$ (10)
@@
$\frac{k^{2}\cdot C_{in}+C_{in}\cdot C_{out}+k^{2}C_{out}}{k^{2}C_{in}\cdot C_{out}+k^{2}C_{out}\cdot C_{out}}=\frac{1}{2C_{in}}+\frac{1}{27}.$ (11)
Observably, compared to the pre-improvement phase, the proposed light-weight residual block diminishes computational and parametric quantities by approximately 27 times when $C_{in}$ is large enough. The exact reduction ratio depends on the $C_{in}$. This novel lightweight residual block, through carefully designed convolution operations and shortcut connections, achieves a reduction in computational load and parameter count while preserving model performance as much as possible, offering new perspectives for the optimization of deep learning models.
@@ IV-A Datasets
Where N denotes the set of recall levels that represent the exact intervals, this experiment takes $R_{11}=\{\frac{1}{10},\frac{2}{10},\frac{3}{10},\cdots,1\}$. $f_{inter}(n)$ denotes the precision obtained after interpolation calculation, instead of averaging the precision values, the maximum precision greater than or equal to the current recall is taken in calculating the precision. It is defined as the Equation (13):
@@
Methods(IoU) Parameters(M) Power(pJ) 3D Object Detection Birds’ Eye View Runtime(s)
@@
ANN SMOKE(0.7) 19.51 2.17E+11 14.76 12.85 11.5 19.99 15.61 15.28 0.03
SMOKE*(0.5) 19.51 2.17E+11 29.16 25.22 21.47 32.52 29.59 25.86 0.03
SMOKE*(0.7) 19.51 2.17E+11 12.03 11.14 10.92 17.97 13.08 12.06 0.03
FCOS3D(0.7) 50.62 3.04E+11 13.9 11.61 10.98 18.16 14.02 13.85
MonoDLE(0.7) 20.82 2.87E+11 17.45 13.66 11.68 24.97 19.33 17.01 0.04
MonoDTR(0.7) 54.25 5.42E+11 24.52 18.57 15.51 33.33 25.35 21.68 0.037
MonoLSS(0.7) 21.51 2.81E+11 25.91 18.29 15.94 34.70 25.36 21.84 0.035
SNN SpikeSMOKE(0.7) 19.51 5.97E+10 8.96 7.49 7.31 12.07 9.95 9.32 0.09
SpikeSMOKE-CSGC(0.7) 19.56 6.04E+10 11.78 10.69 10.48 15.67 13.68 11.83 0.105
SpikeSMOKE(0.5) 19.51 5.97E+10 20.87 19.72 16.89 27.64 22.47 21.65 0.09
SpikeSMOKE-CSGC(0.5) 19.56 6.04E+10 28.83 22.75 19.44 36.23 26.88 25.75 0.105
@@ Methods
NuScenes-mini.
The NuScenes-mini dataset is a subset of the NuScenes dataset, which is collected by 6 cameras, 1 Lidar, 5 millimetre wave radars, GPS and IMU, covering complex traffic conditions as well as variations in weather and light. The dataset contains 8 attributes of 3 types of objects, namely vehicles, bicycles, and pedestrians, and 10 scenes with continuous images of about 20 seconds each, which is an important dataset for 3D object detection in the field of autonomous driving. The dataset presents a composite metric, the nuScenes Detection Score (NDS), which was combined by calculating weighted sums (mAP, mATE, mASE, mAOE, mAVE, and mAAE).
@@ IV-B Experiments Setup
We used random horizontal flip, random scale and shift as data augmentation aiming to increase the diversity of training data. In the network design, we set the number of groups for group normalization to 32 and for less than 32, we take 16. based on the analysis in literature [41], we set $\begin{bmatrix}\bar{h,}\bar{w,}\overline{l,}\end{bmatrix}^{\top}=[1.63,1.53,3.88]^{\top}and\begin{bmatrix}\mu_{z},\sigma_{z}\end{bmatrix}^{\top}=[28.01{\operatorname*{,16.32}}]^{\top}.$ The input image resolution size of the network is 1280x384, and after multiple downsampling, the size is reduced to 32 times of the original size. We trained 172 epoch using four NVIDIA GeForce RTX 4090 GPU, with a size of 4 per batch and a learning rate set to $1.25\times 10^{-4}$. The decay strategy for the learning rate was to reduce the learning rate to 10 times of the original at epoch 47 and 90. We used a threshold of 0.25 to filter the detection object. The lightweight strategy adopted in the experiment only acts on the backbone of feature extraction. Additionally, to more clearly present the configuration of each layer in the network, we have used a tabular format for detailed explanation, as shown in Table VIII. The table lists key information including the size of the convolutional kernels, the number of input and output channels, the type of neurons, and the dimensions of the input and output, all of which are essential parameters when constructing a deep learning model.
TABLE VI: Results on the validation set of the NuScenes-mini dataset.
Methods mAP NDS
SMOKE 31.3 37.2
SpikeSMOKE 25.3 29.8
SpikeSMOKE-CSGC 26.6 31.2
@@
TABLE VII: Ablation experiments on the threshold of spike neurons.
@@
TABLE VIII: The detailed settings of the SpikeSMOKE-CSGC architecture.
Layer Name Input Size Layer Output Size
Base_layer (T,B,3,384,1280) 7×7Conv,32,stride1,padding3→GN→LIF (T,B,32,384,1280)
CSGC (T,B,32,384,1280) Channel-Attn + Spatial-Attn (3×3/5×5/7×7Conv) + Sigmoid (T,B,32,384,1280)
Avgpooling (T,B,32,384,1280) AvgPool2d(kernel=2, stride=2) (T,B,32,192,640)
Layer0 (T,B,32,192,640) 3×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,64,96,320)
Layer1 (T,B,64,96,320) 4×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,128,48,160)
Layer2 (T,B,128,48,160) 6×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,256,24,80)
Backbone Layer3 (T,B,256,24,80) 4×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,512,12,40)
Upblock0 LIF→3×3Conv(512-256)→LIF→ConvTranspose(2×)
Upblock1 LIF→3×3Conv(256-128)→LIF→ConvTranspose(2×)
Neck Upblock2 LIF→3×3Conv(128-64)→LIF→ConvTranspose(2×) (T,B,64,96,320)
class head (T,B,64,96,320) LIF→Conv3×3(64→256)→GN→LIF→Conv1×1(256→3) (B,3,96,320)
Head regression head (T,B,64,96,320) LIF→Conv3×3(64→256)→GN→LIF→Conv1×1(256→8) (B,8,96,320)
@@ IV-C Main Results
We validate the performance on the validation and test sets of the KITTI dataset for 3D object detection and bird’s-eye view under three levels of Easy, Moderate and Hard. * denotes the experimental results obtained by us after configuring the environment and training based on the SMOKE. The 3d object detection of our proposed SpikeSMOKE with CSGC (SpikeSMOKE-CSGC) can reach 28.83/11.78 (Easy), 22.75/10.69 (Moderate), 19.44/10.48 (Hard), and for the BEV detection, they can reach 36.23/15.67 (Easy), 26.88/13.68 (Moderate), 25.75/11.83 (Hard), as shown in Table I, where the metrics is evaluated by $AP|_{R_{11}}$ at 0.5/0.7 IoU threshold. From the results of these experiments, it can be found that although our detection performance has a little gap compared with the SMOKE-ANN*, energy consumption has been significantly reduced. For example, we calculate the energy consumption of the hard category for the 0.7 IoU threshold and find that it can be reduced by 72.2%, while the detection performance is reduced by only 4%. These experimental results are obtained because SMOKE-ANN* relies on continuous value activation and can precisely model 3D geometric attributes. However, SpikeSMOKE converts features into binary spike sequences (0/1), and information is inevitably lost during the encoding process. Meanwhile, the low-power consumption feature of our SNN event-driven mechanism reduces energy efficiency. The formula for calculating energy consumption is as follows:
[…]
where 0.9 denotes the consumption per accumulator operation and 4.6 denotes the multiplier word operation [54][13]. With a time step T=8, we achieved a runtime latency of 0.105s when performing inference on 4 NVIDIA GeForce RTX 4090 GPUs. Since we extended the time dimension when training on the GPU, it is reasonable that the latency would be larger compared to the traditional ANNs model.
@@
We validate the classification results of the proposed CSGC coding strategy on MS-ResNet18 on the classification dataset CIFAR-10/100, as shown in Table IV and Table V. We note that the accuracy rate of the MS-ResNet18 with CSGC coding on the CIFAR-10 is 1.06% higher than the direct coding method. On the CIFAR-100, the proposed method MS-ResNet18-SNN with CSGC coding can achieve 79.58% by 6 time steps, which is 3.17% higher than MS-ResNet18-SNN. Meanwhile, we validated its effectiveness on the NuScenes-mini dataset, as shown in Table VI. Consequently, our proposed CSGC coding method is effective and generalized.
@@
Since the threshold of the spike neurons has a great impact on the detection performance of the model, we conduct some ablation experiments on 2D detection by different neuron thresholds in order to get the most suitable LIF neuron threshold. Based on the experiment results, we obtain that the best performance is obtained when the neuron threshold vth=0.75, which is used in this paper, as shown in Table VII.
@@
We provide visualization results of the monocular 3D object detection on the KITTI dataset and compare them with the SpikeSMOKE and SMOKE, as shown in Figure 5. It can be observed that the detection performance of the Spikesmoke-CSGC architecture is superior to that of the SpikeSMOKE architecture and is also closer to that of the SMOKE architecture. Through these results, we can visually understand the model’s capability to accurately identify and locate 3D objects in complex road scenarios, thereby affirming the significant potential application of this technology in areas such as autonomous driving.
@@ V Conclusion
With the widespread use of 3D object detection in applications such as autonomous driving, the low energy consumption problem is getting more and more attention. As widely acknowledged, low-power consumption stands out as a key feature of brain-like SNNs, offering a potential new solution for energy-efficient 3D object detection. Based on the SMOKE architecture and SNNs model, we have constructed a novel Spiking SMOKE architecture named SpikeSMOKE for monocular 3D object detection to reduce its energy consumption. In this architecture, the discrete signaling characteristics of SNNs may result in information loss and restrict their capacity for feature representation. Therefore, we have proposed a new cross-scale gating coding module named CSGC that can enhance the feature expression ability of the model, inspired by the synaptic filtering process in biological neurons. Furthermore, we have also proposed a lightweight residual block to reduce the computational effort while maintaining the impulse computation paradigm. We have conducted a large number of experiments on the KITTI, NuScenes-mini and CIFAR10/100 datasets to verify that the SpikeSMOKE architecture and CSGC module proposed in this paper are effective. The experimental results on the KITTI dataset, show that SpikeSMOKE can achieve higher energy efficiency compared to SMOKE, e.g., 72.2% reduction in energy consumption on Hard category, while the detection performance drops by only 4%. Moreover, the experimental results also show that the CSGC-based SpikeSMOKE can achieve significant improvement over the baseline SpikeSMOKE. SpikeSMOKE-L (lightweight) can further reduce the amount of parameters by 3 times and computation by 10 times compared to SMOKE. In the CIFAR-10/100 classification task, the CSGC encoding strategy improves the correctness by 1.06% and 3.17%, respectively, validating its generality. Overall, the SpikeSMOKE architecture and CSGC mechanism can provide an efficient and feasible solution for low-power monocular 3D object detection. In the future, we will continue to explore how to utilize the SNNs model to further enhance the performance of monocular 3D object detection and further reduce its energy consumption.